OpenStudy (anonymous):

9.03] Part 1 − Find the vertex, axis of symmetry, domain, and range of the graph of y = −4x2 + 4x − 7. Show all work for full credit. Part 2 − Using complete sentences, explain how you can determine the axis of symmetry, the domain, and range without graphing y = −4x2 + 4x − 7.

OpenStudy (helder_edwin):

do u know how to complete squares?

OpenStudy (anonymous):

no

OpenStudy (helder_edwin):

let's try it to together then: u have \[ \large y=-4x^2+4x-7 \] then \[ \large y+7=-4(x^2-x+\qquad) \] ok so far?

OpenStudy (anonymous):

no why did you put x^2 and -x

OpenStudy (helder_edwin):

first i added +7 to both sides. then i factored -4 from what was left on the right hand side ok?

OpenStudy (anonymous):

ohhh ok

OpenStudy (helder_edwin):

now the coeficent of x^2 is 1 and the coefficient of x is -1. in order to get a perfect square we have to put a number in the blank i left. this number is the coefficient of x halved and then squared. do u get this?

OpenStudy (anonymous):

ok, yes i get it

OpenStudy (helder_edwin):

now we write: \[ \large y+7\color{red}{-4\cdot\left(\frac{-1}{2}\right)^2} =-4\left(x^2-x+\color{red}{\left(\frac{-1}{2}\right)^2}\right) \]

OpenStudy (helder_edwin):

got it?

OpenStudy (helder_edwin):

hello?? @valentinapaez

OpenStudy (helder_edwin):

well u finalli get \[ \large y+6=-4(x-1/2)^2 \]

OpenStudy (helder_edwin):

i hope i was of help. sorry. but i gotta go

OpenStudy (anonymous):

wait